# Generative Enzyme Design Guided by Functionally Important Sites and Small-Molecule Substrates

### v3

@@ D.1 EnzyGen Performance on 323 Testing Fourth-Level Categories
The ESP scores, substrate binding affinities and AlphaFold2 pLDDT results for the 323 testing fourth-level categories are respectively provided in Table 79 and 11.
@@ D.2.1 Top-1 Candidate Results
ESP scores of top-1 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 8,  10 and  12. The tables show that our EnzyGen achieves the best average scores across different categories on all metrics. Notably, the ESP scores of different models consistently surpass those of the top-5 candidate. This observation is reasonable, as candidates with higher probabilities are more likely to achieve higher function scores.
Protein Family 1.1.1 1.11.1 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
@@ D.2.2 Top-5 Candidate Results
ESP scores of top-5 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 13,  14 and  15. The tables show that our EnzyGen achieves the best average scores across different categories on all metrics.
Protein Family 1.1.1 1.11.1 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
@@ D.2.3 Top-10 Candidate Results
Protein Family 1.1.1 1.11.1 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
PROTSEED -4.65 -4.93 -4.76 -11.09 -7.28 -9.46 -$1$9.21 -10.29 -7.88 -5.13 -8.38 -10.66 -11.37 -13.20
@@
ESP scores of top-10 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 16,  17 and  18. Again, our model achieves the best average performance in all cases. It is worth noting that our EnzyGen achieves an average enzyme-substrate ESP score of 0.93, an average substrate binding affinity of -10.12 and an average pLDDT of 90.52, demonstrating EnzyGen’s ability to design both well-folded and functional enzymes.

### v4

@@ D.1 EnzyGen Performance on 323 Testing Fourth-Level Categories
The ESP scores, substrate binding affinities and AlphaFold2 pLDDT results for the 323 testing fourth-level categories are respectively provided in Table 810 and 12.
@@ D.2.1 Top-1 Candidate Results
ESP scores of top-1 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 9,  11 and  13. The tables show that our EnzyGen achieves the best average scores across different categories on all metrics. Notably, the ESP scores of different models consistently surpass those of the top-5 candidate. This observation is reasonable, as candidates with higher probabilities are more likely to achieve higher function scores.
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
@@ D.2.2 Top-5 Candidate Results
ESP scores of top-5 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 14,  15 and  16. The tables show that our EnzyGen achieves the best average scores across different categories on all metrics.
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
@@ D.2.3 Top-10 Candidate Results
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4 Avg
PROTSEED -4.65 -4.93 -4.76 -11.09 -7.28 -9.46 -11.11 -9.21 -10.29 -7.88 -5.13 -8.38 -10.66 -11.37 -13.20
@@
ESP scores of top-10 candidate and their corresponding substrate binding affinities as well as pLDDT are respectively reported in Table 17,  18 and  19. Again, our model achieves the best average performance in all cases. It is worth noting that our EnzyGen achieves an average enzyme-substrate ESP score of 0.93, an average substrate binding affinity of -10.12 and an average pLDDT of 90.52, demonstrating EnzyGen’s ability to design both well-folded and functional enzymes.